the bob of a pendulum is released from horizontal posn a.if the lengthof pendulum is 1.5 m what is the speed with which the bob arrives at lowermost point b, given that it dissipated 5% of its energy against air resistance?


 

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear manisha

initial energy of bob = mg*1.5

energy dissipated = 5/100 * (mg*1.5)


now energy balane

 mg*1.5 = 1/2 m v2 + 5/100 * (mg*1.5)

 v2 =3g - 3g/20

v2 =57g/20

v= √(57g/20)

   
Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin